1. What is the concentration (in M) of an aqueous solution of sulfuric acid prepared by transferring 25.0 mL of 6.00 M H2So4(aq) to a 250 mL-volume volumetric flask that is subsequently filled to its 250 mL-mark with distilled water? 2. What volume of a saturated solution of 8.43 M KI(aq) at 20 C is needed to form 10.0 L of a 1.00 M solution? What volume of pure water is needed as diluent? 3. Identify the type of electrolyte (strong, weak, or non- formed when the following compounds dissolve in pure water. a) Ethylamine (CH3CH2NH2, an organic base) b) Calcium fluoride c) Argon d) oxygen e sodium hydroxide Hexanoic acid 4. Which one of the following solutions is the best conductor of electricity? A. 1 M NaCl B. 1 M NH C. 1 M CH30H D. 1 M CH3COOH E. They all conduct electricity equally well​need help for these questions thank you
Answer #1

Volume of H2SO4 intially V1 = 25 mL = 0.025 L

Molarity of  H2SO4 intially M1 = 6 M (or mol/L)

No. of moles of H2SO4 = V1* M1 = 0.025 L * 6 mol/L = 0.15 moles of H2SO4

same no. of moles of H2SO4 will be present after adding Distilled water to make it to 250 mL = 0.25 L

so Molarity = No. of moles / volume in L = 0.15 moles / 0.25 L = 0.6 mol/L or 0.6 M Answer

Given we need 10 L of 1 M of KCl

so no of moles need to be present = 10 L * 1 mol/L = 10 moles

we have solution of 8.34 M(or mol/L)

so volume of KCl of orginal solution required = No. of moles /Molarity of orginal solution = 10 mol / 8.34 mol/L = 1.12 L

so we require 1.12 L of KCl answer